OpenStudy (anonymous):

A combined area of the countries A and B is 164,973 square kilometers. Country A is larger by 473 square kilometers. Determine the land area of each country.

OpenStudy (anonymous):

164,973-473=164500 164500/2=82250 A: 82250+473=82723 B: 82250